OpenStudy (anonymous):

A recipe calls for 1/3 of a cup of milk for 12 cookies. How many cups of milk are needed to make 72 cookies? 2 24 12 3

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (studygurl14):

Hello @Yaaassss Use proportions. \(\huge\frac{\frac{1}{3}}{12}=\frac{m}{72}\) solve for m.

OpenStudy (studygurl14):

@Yaaaassss